OpenStudy (anonymous):

How would I set this problem up? (Advanced Algebra with Financial Applications) Sean currently has an account balance of $3,981.04. He opened the account five years ago with a deposit of $3,760.19. If the interest compounds daily, what is the interest rate on the account?

OpenStudy (anonymous):

I'm sure once I see the correct formula, I could solve it from there, I just need to know which formula to use. I want to use P(1+r/n)^nt but not exactly sure if that is the correct formula

hero (hero):

\[FV = PV(1 + r)^t\] FV = Future Value PV = Present Value r = interest rate t = total number of times compounded You're given: FV = 3,981.04 PV = 3760.19 t = 365*5

OpenStudy (anonymous):

Okay, thanks to both you! I'll let you know if I am stuck (:

OpenStudy (anonymous):

Okay, so my answer choices are: 1.1% 0.6% 4.6% 5.6% Is 1.1% correct? @Hero

hero (hero):

Correct.

OpenStudy (anonymous):

Sweet thanks!